OpenStudy (anonymous):

f(x) = { x + 5}/ { x + 10 } f^{-1}( -3 ) =? brief steps would be wonderful!

OpenStudy (anonymous):

\[x=\frac{y+5}{y+10}\] sovle for y

OpenStudy (anonymous):

\[x(y+10)=y+5\] \[xy+10x=y+5\] \[xy-y=5-10x\] \[y(x-1)=5-10x\] \[y=\frac{5-10x}{x-1}\]

OpenStudy (anonymous):

They really need another way to write inverse function than f(x)^-1. I get the two confused all the time. >.> You know, like there's arcsin for sin(x)^-1

OpenStudy (anonymous):

oh you want \[f^{-1}(-3)\] so now you have a choice. replace x by -3 in the above solution, or start with \[\frac{x+5}{x+10}=-3\] and solve for x. steps will be the same, except you will get a number rather than an expresson in y

OpenStudy (anonymous):

\[x+5=-3(x+10)\] \[x+5=-3x-30\] \[4x=-35\] \[x=-\frac{35}{4}\] which is just what you would get writing \[f^{-1}(-3)=\frac{5-10(-3)}{-3-1}\]

OpenStudy (anonymous):

thank u really helped simplify it so i could understand it